OpenStudy (anonymous):

points a2,3 and b5,9 round to the nearest tenth

OpenStudy (anonymous):

what do you mean... like are you trying to find the midpoint or what?

OpenStudy (anonymous):

i am guessing it is "find the distance between the points \((2,3)\) and \((5,9)\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

from 2 to 5 in the x direction is right 3 units from 3 to 9 in the y is up 6 pythagoras gives the distance as \[d^2=3^2+6^2=9+36=45\] and so \[d=\sqrt{45}\]

OpenStudy (anonymous):

you are welcome to use the distance formula if you like \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] but the idea is pythagoras